Hello, I am having problems solving this equation. Please assist me in learning. :)
7 - 8r + 3 = 6r - 12 + r
the lone r on the end, is that 1r and do I add 1r to 6r?
I'm confused what do do with that r thats alone
Thank you, Steve

Answer by stanbon(75887)

7 - 8r + 3 = 6r - 12 + r
the lone r on the end, is that 1r and do I add 1r to 6r?
---
Yes; just think of "r" are cars.
Then you have
7 - 8cars + 3 = 6cars - 12 + 1car
-----
7 -8 cars + 3 = 7cars -12
---
10-8cars = 7cars -12
------
15cars = 22
cars = 22/15
----
r = 22/15
=============
Cheers,
Stan H.